# Ensemble Kalman Filtering Meets Gaussian Process SSM for Non-Mean-Field and Online Inference

### v3

@@ IV-A Ensemble Kalman Filter (EnKF)
The EnKF is a Monte Carlo-based method that excels in handling nonlinear systems compared to the KF and EKF [40]. Given a SSM, see Eq. (II-B), the EnKF sequentially approximates the filtering distributions using $N\!\in\!\mathbb{N}$ equally weighted particles [40]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
With the linear Gaussian emission in Eq. (II-B), the joint distribution of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ can be readily obtained as follows:
@@
Compared to another Monte Carlo-based method, PF, EnKF demonstrates computational efficiency, particularly in exceedingly high-dimensional spaces [41], though the PF may offer more flexibility in addressing highly non-Gaussian and nonlinear aspects. More importantly, the inherent differentiable nature of the EnKF (see Eqs. (10)–(14)) enables seamless integration with off-the-shelf automatic differentiation tools (e.g., PyTorch [42]) for model parameter optimization (see more discussions in Section IV-C). This stands in contrast to PF, where the discrete distribution resampling significantly poses challenges for the utilization of the reparameterization trick [43], resulting in significantly higher computational complexity for the parameters gradient computations [44, 45, 46, 47].
In the following subsections, we will describe our novel EnKF-aided NMF variational inference method, which significantly reduces the number of variational parameters and ultimately enhances the learning and inference performance of GPSSMs.
@@ IV-B1 Sparse GPSSMs
where $p({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})p({\vec{\mathbf{f}}}|{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the augmented GP prior and $p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the noiseless GP prediction whose mean and covariance can be computed similarly to Eq. (II-A). The introduced inducing inputs ${\vec{\mathbf{z}}}$ will be treated as variational parameters and jointly optimized with model parameters [48]. We will further describe this later.
@@ IV-B2 ELBO for sparse GPSSM
where $\mathbf{m}\!\!=\!\![{\mathbf{m}}_{1}^{\top},\ldots,{\mathbf{m}}_{d_{x}}^{\top}]^{\top}\in\mathbb{R}^{Md_{x}}$ and $\mathbf{S}\!\!=\!\!\operatorname{diag}(\mathbf{L}_{1}\mathbf{L}_{1}^{\top},\ldots,\mathbf{L}_{d_{x}}\mathbf{L}_{d_{x}}^{\top})\in\mathbb{R}^{Md_{x}\times Md_{x}}$ are free variational parameters. This explicit representation of the variational distribution enables scalability through the utilization of stochastic gradient-based optimization [49], as it allows for the independence of individual GP predictions given the explicit inducing points [48]. Similarly, we assume that the variational distribution over the initial state is $q({\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{0}|{\mathbf{m}}_{0},\mathbf{L}_{0}\mathbf{L}_{0}^{\top})$, where ${{\mathbf{m}}}_{0}\in\mathbb{R}^{d_{x}}$ and lower-triangular matrix $\mathbf{L}_{0}\in\mathbb{R}^{d_{x}\times d_{x}}$ are free variational parameters of $q({\mathbf{x}}_{0})$. The variational distribution of the latent states, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, obtained by integrating out ${\mathbf{f}}_{t}$ as shown in Eq. (19), will be implicitly modeled by resorting to the EnKF technique. This modeling approach will help eliminate the need to parameterize $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, a common requirement in the existing works [26, 27, 28, 29], thus overcoming the challenges associated with optimizing a large number of variational parameters, as discussed in Section III.
@@
Upon the augmentation of induced points into the GPSSM (see Eq. (16)) and under the NMF assumption of the variational distribution (see Eq. (19)), the model evidence lower bound for joint learning and inference is:
@@
An important observation lies in the interpretability of each component within the ELBO in Eq. (21). Specifically, when maximizing the ELBO for learning and inference:
@@ IV-B3 EnVI algorithm
The NMF assumption applied to the variational distribution (Eq. (19)) typically results in an intractable evaluation of the first and fourth terms within the ELBO [28]. We next proceed to outline the approximate evaluation of the ELBO using EnKF. The main idea revolves around utilizing the filtering distribution acquired from EnKF as a proxy for the corresponding variational distribution, acting as an approximation for the underlying posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
Building upon the EnKF outlined in Section IV-A and assuming that we have acquired the variational distribution, $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\!=\!{\cal N}({\mathbf{x}}_{t-1}|\mathbf{m}_{t-1},\mathbf{P}_{t-1})$, at time step $t\!-\!1$, we can employ the GP transition, presented in Eq. (17), to perform the prediction step and generate $N$ predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, i.e.,
where $\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$ are $N$ particles generated from $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ with equal weights. With the predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, we thus can approximate the prediction distribution as a Gaussian:
$=\int p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\mathrm{d}{\mathbf{x}}_{t-1}$
$\approx{\cal N}({\mathbf{x}}_{t}|\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t})$ (23)
where $\bar{\mathbf{m}}_{t}$ and $\bar{\mathbf{P}}_{t}$ can be computed using Eq. (IV-A). Similarly, during the update step, utilizing Eqs. (IV-A) and (14), we can derive the variational distribution (approximate posterior distribution) at time step $t$,
$q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})\triangleq p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})={\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t})$ (24)
and obtain the set of $N$ updated samples ${\mathbf{x}}_{t}^{1:N}$. Note that in the context of GPSSMs, both the prediction and update steps are conditioned on the inducing points, which act as a surrogate for the transition function.
So far, we have introduced the utilization of EnKF to obtain the approximate posterior distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$. However, in order to evaluate the ELBO and compute term 1 and term 4 in Eq. (21), we still need to determine the conditional distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Despite having that
$q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})=\int q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})\mathrm{d}{\mathbf{x}}_{t-1},$ (25)
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to simply let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [27], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational lower bound approximation. The main result is summarized in the following proposition.
@@
Under the approximations of EnKF filtering (Eq. (24)) and $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$, the ELBO presented in Theorem 1 can be reformulated as a summation over several simple terms:
@@
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in the first term can be analytically evaluated, i.e.,
@@
and the two KL divergence terms can also be analytically computed in closed form, due to the Gaussian nature of the prior and variational distributions [10].
Proof.
The proof can be found in Appendix A
Based on the result in Proposition 1, we first utilize the reparametrization trick [43] to sample ${\vec{\mathbf{u}}}\!\sim\!q({\vec{\mathbf{u}}})$ and numerically get an unbiased evaluation of the expected log-likelihood, $\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$; then we apply gradient-based methods [50] to maximize the $\mathcal{L}$ in Eq. (26) w.r.t. the model parameters ${{\bm{\theta}}}=[\bm{\theta}_{gp},\mathbf{Q},\mathbf{R}]$ and the variational parameters $\bm{\zeta}=[\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}]$. Detailed routine for implementing the EnKF-aided variational learning and inference algorithm, termed as EnVI, is summarized in Algorithm 1.
@@
1:Input: ${{\bm{\theta}}}=[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}],~{}{\bm{\zeta}},~{}{\mathbf{y}}_{1:T}$, ${\mathbf{x}}_{0}^{1:N}\sim q(x_{0})$

### v5

@@ IV-A Ensemble Kalman Filter (EnKF)
The EnKF is a Monte Carlo-based method that excels in handling nonlinear systems compared to the KF and EKF [45]. Given an SSM, see Eq. (4), the EnKF sequentially approximates the filtering distributions using $N\!\in\!\mathbb{N}$ equally weighted particles [45]. Specifically, at the prediction step, EnKF first samples particles, ${\mathbf{x}}_{t-1}^{1:N}\triangleq\{{\mathbf{x}}_{t-1}^{n}\}_{n=1}^{N}$, from the filtering distribution, $p({\mathbf{x}}_{t-1}|{\mathbf{y}}_{1:t-1})$, at time $t-1$. Then, the sampled particles are propagated using the state transition function $f(\cdot)$, i.e.,
@@
With the linear Gaussian emission in Eq. (4), the joint distribution of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ can be readily obtained as follows:
@@
It is crucial to note that the EnKF inherently exhibits differentiability. Specifically, if we utilize the reparameterization trick [46] to sample the process and observation noises as shown in Eq. (10) and (14), i.e.,
$\mathbf{v}_{t}^{n}=\bm{0}+\mathbf{Q}^{\frac{1}{2}}\bm{\epsilon},\ \ \ \bm{\epsilon}\sim{\cal N}(\bm{0},\bm{I}_{d_{x}}),$ (15a)
$\mathbf{e}_{t}^{n}=\bm{0}+\mathbf{R}^{\frac{1}{2}}\bm{\epsilon},\ \ \ \bm{\epsilon}\sim{\cal N}(\bm{0},\bm{I}_{d_{y}}),$ (15b)
then the sampled latent states become differentiable w.r.t. the state transition function $f$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This differentiable nature is crucial for enabling the use of gradient-based optimization methods in learning algorithms. Moreover, although we present the linear emission model in this paper, the EnKF can be extended to nonlinear emission models. The difference lies in the computation of the predicted observation covariance and the cross-covariance between the state and the observation predictions in Eq. (12), which are done using ensembles. For more details, see e.g. [45].
@@
Compared to another Monte Carlo-based method, PF, EnKF demonstrates computational efficiency, particularly in exceedingly high-dimensional spaces [47], though the PF may offer more flexibility in addressing highly non-Gaussian and nonlinear aspects. Moreover, the inherent differentiable nature of EnKF (see Eqs. (10)–(14)) enables seamless integration with off-the-shelf automatic differentiation tools (e.g., PyTorch [48]) for model parameter optimization (see more discussions in Section IV-C). This stands in contrast to PF, where the discrete distribution resampling significantly poses challenges for the utilization of the reparameterization trick [46], resulting in significantly higher computational complexity for the parameters gradient computations [49, 50, 51].
In the following subsection, we will describe our novel EnKF-aided NMF variational inference algorithm for GPSSMs, which significantly reduces the number of variational parameters and ultimately enhances the learning and inference performance.
@@ IV-B1 Sparse GPSSMs
where $p({\vec{\mathbf{f}}},{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})p({\vec{\mathbf{f}}}|{\vec{\mathbf{u}}})=p({\vec{\mathbf{u}}})\prod_{t=1}^{T}p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the augmented GP prior and $p({\mathbf{f}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}})$ is the noiseless GP prediction whose mean and covariance can be computed similarly to Eq. (3). The introduced inducing inputs ${\vec{\mathbf{z}}}$ will be treated as variational parameters and jointly optimized with model parameters [52]. We will further describe this later.
@@ IV-B2 ELBO for sparse GPSSM
where $\mathbf{m}\!\!=\!\![{\mathbf{m}}_{1}^{\top},\ldots,{\mathbf{m}}_{d_{x}}^{\top}]^{\top}\in\mathbb{R}^{Md_{x}}$ and $\mathbf{S}\!\!=\!\!\operatorname{diag}(\mathbf{L}_{1}\mathbf{L}_{1}^{\top},\ldots,\mathbf{L}_{d_{x}}\mathbf{L}_{d_{x}}^{\top})\in\mathbb{R}^{Md_{x}\times Md_{x}}$ are free variational parameters. This explicit representation of the variational distribution enables scalability through the utilization of stochastic gradient-based optimization [53], as it allows for the independence of individual GP predictions given the explicit inducing points [52]. The corresponding inducing inputs, ${\vec{\mathbf{z}}}$, are treated as variational parameters as well, as described in [34, 52]. These parameters, $\{{\vec{\mathbf{z}}},{\mathbf{m}},\mathbf{S}\}$, collectively define the variational distribution that approximates the true GP posterior. We also assume that the variational distribution over the initial state is $q({\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{0}|{\mathbf{m}}_{0},\mathbf{L}_{0}\mathbf{L}_{0}^{\top})$, where ${{\mathbf{m}}}_{0}\in\mathbb{R}^{d_{x}}$ and lower-triangular matrix $\mathbf{L}_{0}\in\mathbb{R}^{d_{x}\times d_{x}}$ are free variational parameters of $q({\mathbf{x}}_{0})$. The variational distribution of the latent states, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, obtained by integrating out ${\mathbf{f}}_{t}$ as shown in Eq. (20), will be implicitly modeled by resorting to the EnKF technique. This modeling approach will help eliminate the need to parameterize $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, a common requirement in the existing works [28, 29, 30, 31], thus overcoming the challenges associated with optimizing a large number of variational parameters, as discussed in Section III.
@@
Upon the augmentation of the inducing points into the GPSSM (see Eq. (17)) and under the NMF assumption of the variational distribution (see Eq. (20)), the model evidence lower bound for joint learning and inference is:
@@
An important observation lies in the interpretability of each component within the ELBO. Specifically, when maximizing the ELBO for learning and inference:
@@ IV-B3 EnVI algorithm
The NMF assumption applied to the variational distribution (Eq. (20)) typically results in an intractable evaluation of the first and fourth terms within the ELBO [30]. Therefore, approximations are needed to help efficiently evaluate the ELBO. One existing method is to simply set $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})\!=\!p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, resulting in the same ELBO as in [29], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (22) and propose our variational lower bound approximation. The main result is summarized in the following proposition.
@@
Under the approximations that:
  • 1)
    $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\approx p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$,
  • 2)
    $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})\approx p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$,
the ELBO presented in Theorem 1 can be reformulated as a summation over several simple terms:
@@
where the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in the first term can be analytically evaluated using the EnKF (discussed below). The two KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [17].
Proof.
The proof can be found in Appendix A. ∎
Proposition 1 demonstrates that the newly derived ELBO is significantly more tractable under two mild approximations, which are justified and explained as follows.
  • Approximation 1) posits that the state estimation at time $t\!\!-\!\!1$ using observations from time $1$ to $t\!\!-\!\!1$ (i.e., $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$) is approximately equal to the estimation using observations from time $1$ to $t$ (i.e., $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})$). This approximation is generally reasonable, particularly when $t$ is large (i.e., with a long observation sequence), as the increased information aids in more accurately inferring the latent state, thereby reducing the discrepancy between the two posterior distributions.
  • Approximation 2) states that the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$ is approximately equal to the posterior distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$. In the variational inference framework [53], the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$ is intended to approximate the true but unknown distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:T})$. That is to say, in Approximation 2), we are essentially using the filtering distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:t})$ to approximate the smoothing distribution $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1},{\mathbf{y}}_{1:T})$. While this approximation may introduce some estimation loss, it remains reasonable when the observation series is sufficiently long, such that additional future observations do not significantly alter the state estimates. On the other hand, given our priority on computational efficiency, a minor loss in accuracy is acceptable and often necessary as part of this trade-off.
We next proceed to outline the evaluation of the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (23) using the EnKF. which allows us to evaluate the log-likelihood recursively and analytically.
Building upon the EnKF outlined in Section IV-A and assuming that we have acquired the posterior distribution, $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\!=\!{\cal N}({\mathbf{x}}_{t-1}|\mathbf{m}_{t-1},\mathbf{P}_{t-1})$, at time $t\!-\!1$, we can employ the GP transition, presented in Eq. (18), to perform the prediction step and generate $N$ predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, i.e.,
where ${\mathbf{x}}_{t-1}^{1:N}$ are $N$ particles obtained from $p({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ with equal weights. With the predicted samples $\bar{{\mathbf{x}}}_{t}^{1:N}$, we thus can approximate the prediction distribution as a Gaussian:
$\approx{\cal N}({\mathbf{x}}_{t}|\bar{\mathbf{m}}_{t},\bar{\mathbf{P}}_{t})$ (25)
where $\bar{\mathbf{m}}_{t}$ and $\bar{\mathbf{P}}_{t}$ can be computed using Eq. (11).
Similarly, during the update step, utilizing Eqs. (13) and (14), we can derive the filtering distribution at time step $t$,
$p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t})={\cal N}({\mathbf{x}}_{t}|\mathbf{m}_{t},\mathbf{P}_{t})$ (26)
and obtain the set of $N$ updated samples ${\mathbf{x}}_{t}^{1:N}$. We can then recursively obtain the samples and posterior distributions in the subsequent time steps. Note that in the context of GPSSMs, both the prediction and update steps are conditioned on the inducing points, $\{{\vec{\mathbf{u}}},{\vec{\mathbf{z}}}\}$, which act as a surrogate for the transition function.
Leveraging the EnKF outlined above, the log-likelihood, $\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})$ in Eq. (23), can be evaluated recursively and analytically. Specifically, at each time step $t$, we have
@@
due to the Gaussian prediction distribution, see Eq. (25), and the linear emission model.
Now we can evaluate our approximate variational lower bound, $\mathcal{L}$ in Eq. (23). We first utilize the reparameterization trick (see also e.g. Eq. (15)) [46] to sample ${\vec{\mathbf{u}}}$, i.e.,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\vec{\mathbf{u}}}=\mathbf{m}+\mathbf{S}^{\frac{1}{2}}\bm{\epsilon},\ \bm{\epsilon}\sim{\cal N}(\bm{0},\bm{I}),}$ (28)
and numerically get an unbiased evaluation of the expected log-likelihood, $\mathbb{E}_{q({\vec{\mathbf{u}}})}\!\left[\sum_{t=1}^{T}\log p({\mathbf{y}}_{t}|{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t-1})\right]$. Due to the reparameterization trick [46], $\mathcal{L}$ is differentiable w.r.t. the model parameters ${{\bm{\theta}}}\!=\!\{\bm{\theta}_{gp},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m},\mathbf{S},{\vec{\mathbf{z}}}\}$. Therefore, we can use modern differentiation tools, such as PyTorch, to automatically compute the gradient through backpropagation through time (BPTT) and apply gradient-based methods (e.g., Adam) to maximize $\mathcal{L}$ [48, 54]. Detailed routine for implementing the EnKF-aided variational learning and inference algorithm, termed as EnVI, is summarized in Algorithm 1.
@@
1:Input: ${{\bm{\theta}}}=\{{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}\},~{}{\bm{\zeta}},~{}{\mathbf{y}}_{1:T}$, ${\mathbf{x}}_{0}^{1:N}\sim q(x_{0})$